# Epsilon-Greedy Thompson Sampling to Bayesian Optimization

### v2

@@ Abstract
Bayesian optimization (BO) has become a powerful tool for solving simulation-based engineering optimization problems thanks to its ability to integrate physical and mathematical understandings, consider uncertainty, and address the exploitation–exploration dilemma. Thompson sampling (TS) is a preferred solution for BO to handle the exploitation–exploration trade-off. While it prioritizes exploration by generating and minimizing random sample paths from probabilistic models—a fundamental ingredient of BO—TS weakly manages exploitation by gathering information about the true objective function after it obtains new observations. In this work, we improve the exploitation of TS by incorporating the $\varepsilon$-greedy policy, a well-established selection strategy in reinforcement learning. We first delineate two extremes of TS, namely the generic TS and the sample-average TS. The former promotes exploration, while the latter favors exploitation. We then adopt the $\varepsilon$-greedy policy to randomly switch between these two extremes. Small and large values of $\varepsilon$ govern exploitation and exploration, respectively. By minimizing two benchmark functions and solving an inverse problem of a steel cantilever beam, we empirically show that $\varepsilon$-greedy TS equipped with an appropriate $\varepsilon$ is more robust than its two extremes, matching or outperforming the better of the generic TS and the sample-average TS.
@@ 1 Introduction
where ${\bf x}\in\mathbb{R}^{d}$ is the vector of $d$ input variables selected in a bounded, compact domain $\mathcal{X}\subset\mathbb{R}^{d}$, and $f({\bf x}):\mathcal{X}\mapsto\mathbb{R}$ is a real-valued objective function. In science and engineering applications, the objective function $f({\bf x})$ is often a black-box function evaluated via costly simulations, which unfortunately hinders the use of any mature, derivative-based numerical optimization programs.
@@
Bayesian optimization (BO) [1, 2, 3, 4, 5] is a global sequential optimization technique well-suited for solving small- and moderate-dimensional optimization problems with costly or black-box objective functions. It finds application in diverse domains of science and engineering, including hyperparameter tuning of machine learning algorithms [1], identification of material parameters [6, 7], material design [8, 9], airfoil design [10], adaptive experimental design [11], and accelerator physics [12]. For a recent comprehensive review of its applications in engineering design, see [5]. At its core, BO guides the optimization process using a probabilistic surrogate model $\widehat{f}$ of the objective function, usually a Gaussian process (GP), coupled with an optimization policy [4]. Given a dataset that has several observations of the input variables and the corresponding objective function values, a GP posterior built from this dataset often serves as the probabilistic model $\widehat{f}$ encapsulating our beliefs about the black-box objective function. The optimization policy specifying what we value in the dataset is defined through an acquisition function $\alpha(\mathbf{x})$. This acquisition function can be deterministic or stochastic, and is cost-effective to evaluate given $\widehat{f}$, making it convenient for processing optimization. Different considerations that should be taken into account when deriving $\alpha(\mathbf{x})$ from $\widehat{f}$ include, for example, the value of the objective function [13], the information about the minimum location [14, 15, 16], and the information about the minimum value [17].
@@
Various algorithms have been developed to address the classic exploitation–exploration dilemma in BO, where exploitation involves selecting new solutions anticipated to improve the objective function value and exploration entails sampling new solutions to reduce uncertainty in predictions of the objective function. An effective BO algorithm is deemed to balance these opposing concerns [4]. Notable deterministic acquisition functions such as expected improvement (EI) [18], weighted EI [19], GP upper confidence bound (GP-UCB) [20], knowledge gradient (KG) [21], and likelihood-weighting [22] can manage the exploitation–exploration trade-off. However, they are considered myopic policies, and therefore tend to favor exploitation [13].
@@
Thompson sampling (TS) is a preferred algorithm to address the exploitation–exploration dilemma in solving multi-armed bandit problems [23, 24, 25]. More specifically, TS selects arms (or actions) from a set of possible arms over a limited number of iterations. Each arm corresponds to a stochastic reward drawn from an unknown distribution. The goal is to craft a sequence of arms that maximizes the cumulative reward assuming that the rewards are independent of time and conditionally independent given the selected arms. When applied to GP-based BO, TS generates a sequence of solution points using a mechanism that involves random sampling from an unknown posterior distribution of the global minimum location $\bf{x}^{\star}$ [26]. Thus, we can consider TS a BO method that uses a stochastic acquisition function. It is also worth noting that several information-theoretic optimization policies of BO, such as predictive entropy [16] and max-value entropy [17], compute their acquisition functions based on a set of samples generated by TS.
@@
While TS naturally manages the exploitation–exploration trade-off, its randomness can downplay the role of exploitation in optimization. When the GP posterior exhibits high uncertainty, TS prioritizes exploration as it can diversify the selection of new solutions if there is not much information about $\bf{x}^{\star}$ we have gained. As the number of observations increases, the algorithm transitions to exploiting knowledge about the true objective function. This exploitation strategy, however, is inferior due to the randomness of TS, motivating the quest for an intervention to improve its exploitation.
@@
Several works have explored the $\varepsilon$-greedy policy in BO and multi-armed bandit problems. De Ath et al. [28] proposed two schemes for applying the policy to BO. The first scheme performs exploration (with probability $\varepsilon$) by randomly selecting a point on the Pareto frontier obtained by simultaneously minimizing the posterior mean and maximizing the posterior standard deviation. The second scheme performs exploration by randomly selecting a point in the input variable space. Both schemes implement exploitation (with probability $1-\varepsilon$) by minimizing the posterior mean function. Jin et al. [29] introduced the so-called $\varepsilon$-exploring TS to multi-armed bandit problems. Given the posterior distributions of aims, exploration (with probability $\varepsilon$) sets the estimated reward function for each aim as a random sample drawn from the associated posterior distribution, while exploitation (with probability $1-\varepsilon$) sets the estimated reward function as the sample mean function. Yet the performance of $\varepsilon$-greedy TS for BO remains unexplored.
Our contributions are as follows: (1) introducing $\varepsilon$-greedy TS to BO, which is a simple, effective method to improve the exploitation of TS; and (2) empirical evaluations demonstrating that $\varepsilon$-greedy TS with an appropriate $\varepsilon$ is more robust than its two extremes (i.e., the generic TS and the sample-average TS), matching or outperforming the better of the two across various problems.
@@ 2.1 Overview of Gaussian Processes
Let $\mathcal{D}=\left({\bf X},{\bf y}\right)=\left({\bf x}^{i},y^{i}\right)_{i=1}^{N}$ denote a training dataset, where $\textbf{x}^{i}$ are observations of a $d$-dimensional vector of input variables and $y^{i}$ the corresponding observations of the objective function. We wish to build from $\mathcal{D}$ an observation model $y({\bf x})=f({\bf x})+\varepsilon_{\text{n}}:\mathbb{R}^{d}\mapsto\mathbb{R}$, where $\varepsilon_{\text{n}}\sim\mathcal{N}(0,\sigma^{2}_{\text{n}})$ is additive zero-mean Gaussian noise with variance $\sigma^{2}_{\text{n}}$. This observation noise is assumed to be independent and identically distributed.
A GP model assumes that any finite subset of an infinite set of objective function values has a joint Gaussian distribution [30]. Oftentimes, this assumption is encoded using the following GP prior:
@@
where $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})=\text{cov}[f({\bf x}),f({\bf x}^{\prime})]:\mathbb{R}^{d}\times\mathbb{R}^{d}\mapsto\mathbb{R}$ is a positive semi-definite covariance function parameterized by a vector $\boldsymbol{\theta}$ of hyperparameters. Common-used covariance functions include squared exponential (SE), Matérn 3/2, Matérn 5/2, and automatic relevance determination squared exponential (ARD SE) [4, 30]. By conditioning on $\mathcal{D}$ the model prior given in Eq. 2 and utilizing the observation noise assumption, it is straightforward to show that the vector of observations $\left[y({\bf x}^{1}),\dots,y({\bf x}^{N})\right]^{\intercal}$ is distributed according to an $N$-variate Gaussian with zero mean and covariance matrix ${\bf C}={\bf K}+\sigma^{2}_{\text{n}}{\bf I}_{N}$, where the $(i,j)$th element of ${\bf K}$ is $\kappa({\bf x}^{i},{\bf x}^{j}|\boldsymbol{\theta})$ and ${\bf I}_{N}$ the $N$-by-$N$ identity matrix. By further applying the conditional multivariate Gaussian [31], we obtain the posterior predictive distribution of the objective function value at an unseen input variable vector ${\bf x}_{\star}$, i.e., $p\left(f({\bf x}_{\star})|{\bf x}_{\star},\mathcal{D}\right)=\mathcal{N}\left(\mu_{\text{f}}({\bf x}_{\star}),\sigma_{\text{f}}^{2}({\bf x}_{\star})\right)$, which encodes information about the observation model we wish to build. The mean and variance of the predictive distribution read
@@
Training GP models involves finding an optimal set of hyperparameters $\boldsymbol{\theta}$ that ensures robust predictive performance of trained GPs across unseen input variables. Exact training methods emulating all possible hyperparameter hypotheses are often impractical, leading to popular use of approximate training methods. These approximate training methods fall into two main categories: deterministic approximations and Monte Carlo methods [32]. Deterministic approximations include the maximum likelihood method identifying a set of hyperparameters that maximizes the parameter likelihood $p(\mathcal{D}|\boldsymbol{\theta})$, and Laplace’s method approximating the posterior $p(\boldsymbol{\theta}|\mathcal{D})$ as a multivariate Gaussian with a posterior mean equal to the maximum likelihood estimation and a posterior covariance estimated from Laplace approximation. Monte-Carlo methods, such as importance sampling, the Metropolis, Gibbs sampling, and slice sampling, generate approximate samples of $\boldsymbol{\theta}$ from the posterior $p(\boldsymbol{\theta}|\mathcal{D})$ given the likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ and a prior $p(\boldsymbol{\theta})$. Several of the aforementioned approximate training methods have been incorporated in reliable GP toolboxes such as DACE [33], GPML [34], GPstuff [35], pyGPs [36], and GPflow [37].
@@
1:input: input variable domain $\mathcal{X}$, number of initial observations $N$, threshold for number of BO iterations $K$
@@ 2.2 Bayesian Optimization via Thompson Sampling
Figure 1 shows several sample paths generated from the GP posterior of univariate function $f(x)=x\sin{x}$ for $x\in[0,20]$ and an approximate distribution of their minimum locations. Specifically, we build from ten observations a GP model with zero mean and SE covariance function. From this GP model, we generate 50 sample paths and five of them are shown in Fig. 1(a). Minimizing the generated sample paths and utilizing a kernel density estimation for the obtained solutions result in an approximate distribution $p({\bf x}^{\star}|\mathcal{D})$, as shown in Fig. 1(b). Note that the generic TS in Algorithm 1 generates and minimizes only one sample path in each iteration and does not attempt to approximate the posterior distribution of the minimum location. The following section describes how we can randomly generate sample path $g({\bf x}|\mathcal{D}^{k-1})$ in 10 of Algorithm 1.
Table 1: Spectral density functions associated with common-used covariance functions [38].
Covariance function $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ Spectral density function $p({\bf s})$
@@
2.3 Sampling from Gaussian Process Posteriors
Given the GP posterior $\widehat{f}^{k}({\bf x})|\mathcal{D}^{k-1}$ characterized by a stationary covariance function, we follow the spectral sampling approach by Hernández-Lobato at el. [16] to generate a sample path $g({\bf x}|\mathcal{D}^{k-1})$. This approach approximates the GP prior in Eq. 2 using a Bayesian linear model of randomized basis functions to avoid the computational cost due to exhaustive sampling from marginal distributions of the objective function values at finite sets of input locations, which scales cubically in the number of input locations. Such an approximation is rooted in Bochner’s theorem that guarantees the existence of a Fourier dual $p({\bf s})$ (${\bf s}\in\mathbb{R}^{d}$) of the stationary covariance function, which is called spectral density when a finite non-negative Borel measure is interpreted as a distribution [39]. The spectral density functions associated with several covariance functions of Matérn class and ARD SE are listed in Table 1. Once $p({\bf s})$ is determined, we can represent $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ using the following randomized feature map [40]:
@@ 2.3 Sampling from Gaussian Process Posteriors
where the feature map $\boldsymbol{\phi}({\bf x})\in\mathbb{R}^{N_{\text{p}}}$ can be approximated by [16]
@@
Here, ${\bf W}\in\mathbb{R}^{N_{\text{p}}\times d}$ and ${\bf b}\in\mathbb{R}^{N_{\text{p}}}$ stack $N_{\text{p}}$ spectral points generated from $p({\bf s})$ and $N_{\text{p}}$ points drawn from the uniform distribution $\mathcal{U}[0,2\pi]$, respectively. $\kappa(0|\boldsymbol{\theta})$ is well defined because $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ is a stationary covariance function satisfying $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})=\kappa(\left\|{\bf x}-{\bf x}^{\prime}\right\||\boldsymbol{\theta})$.
Figure 2 shows approximate SE and Matérn 5/2 covariance functions computed from Eq. 6 for different numbers of spectral point samples. We see that, the approximate covariance functions converge to the true ones when increasing the number of spectral points.
where the prior of weight vector $\boldsymbol{\beta}$ is $\mathcal{N}({\bf 0},{\bf I}_{N_{\text{p}}})$. By further conditioning Eq. 8 on the data, we obtain the following mean and covariance matrix of posterior $\boldsymbol{\beta}$ [16]
where $\boldsymbol{\Phi}=[\boldsymbol{\phi}({\bf x}),\dots,\boldsymbol{\phi}({\bf x}^{N})]^{\intercal}\in\mathbb{R}^{N\times N_{\text{p}}}$. To this end, we can approximate the GP posterior using a Bayesian linear model that weights the randomized basis functions associated with the posterior covariance function by samples from $\boldsymbol{\beta}$ posterior.
The generation of $g({\bf x}|\mathcal{D}^{k-1})$ using random features is detailed in Algorithm 2. Note that 10 of Algorithm 2 involves the Cholesky decomposition of matrix $\boldsymbol{\Phi}^{\intercal}\boldsymbol{\Phi}+\sigma_{\text{n}}^{2}{\bf I}_{N_{\text{p}}}$ to facilitate the calculation of its inverse. Alternative posterior sampling techniques such as the decoupled sampling method based on updating GP prior samples [41] can be adopted to increase the sample path accuracy.
@@
1:Input: dataset $\mathcal{D}^{k-1}$, type of stationary covariance function $\kappa(\cdot,\cdot|\boldsymbol{\theta})$, number of spectral points $N_{\text{p}}$

### v3

@@ Abstract
Bayesian optimization (BO) has become a powerful tool for solving simulation-based engineering optimization problems thanks to its ability to integrate physical and mathematical understandings, consider uncertainty, and address the exploitation–exploration dilemma. Thompson sampling (TS) is a preferred solution for BO to handle the exploitation–exploration trade-off. While it prioritizes exploration by generating and minimizing random sample paths from probabilistic models—a fundamental ingredient of BO—TS weakly manages exploitation by gathering information about the true objective function after it obtains new observations. In this work, we improve the exploitation of TS by incorporating the $\varepsilon$-greedy policy, a well-established selection strategy in reinforcement learning. We first delineate two extremes of TS, namely the generic TS and the sample-average TS. The former promotes exploration, while the latter favors exploitation. We then adopt the $\varepsilon$-greedy policy to randomly switch between these two extremes. Small and large values of $\varepsilon$ govern exploitation and exploration, respectively. By minimizing two benchmark functions and solving an inverse problem of a steel cantilever beam, we empirically show that $\varepsilon$-greedy TS equipped with an appropriate $\varepsilon$ is more robust than its two extremes, matching or outperforming the better of the generic TS and the sample-average TS. [^1] This is the accepted version of the following article: Do, B.; Adebiyi, T. & Zhang, R. Epsilon-Greedy Thompson Sampling to Bayesian Optimization. Journal of Computing and Information Science in Engineering, (2024), https://doi.org/10.1115/1.4066858.
@@ 1 Introduction
where ${\bf x}\in\mathbb{R}^{d}$ is the vector of $d$ input variables selected in a bounded, compact domain $\mathcal{X}\subset\mathbb{R}^{d}$, and $f({\bf x}):\mathcal{X}\mapsto\mathbb{R}$ is a real-valued objective function. In science and engineering applications, the objective function $f({\bf x})$ is often a black-box function evaluated via costly simulations, which unfortunately hinders the use of any mature, gradient-based numerical optimization algorithms.
@@
Bayesian optimization (BO) [1, 2, 3, 4, 5, 6] is a global sequential optimization technique well-suited for solving small- and moderate-dimensional optimization problems with costly or black-box objective functions. It finds application in diverse domains of science and engineering, including hyperparameter tuning of machine learning algorithms [1], identification of material parameters [7, 8], material design [9, 10], airfoil design [11], adaptive experimental design [12], and accelerator physics [13]. For a recent comprehensive review of its applications in engineering design, see [5]. At its core, BO guides the optimization process using a probabilistic surrogate model $\widehat{f}$ of the objective function, usually a Gaussian process (GP), coupled with an optimization policy [4]. Given a dataset that has several observations of the input variables and the corresponding objective function values, a GP posterior built from this dataset often serves as the probabilistic model $\widehat{f}$ encapsulating our beliefs about the black-box objective function. The optimization policy specifying what we value in the dataset is defined through an acquisition function $\alpha(\mathbf{x})$. This acquisition function can be deterministic or stochastic, and is cost-effective to evaluate given $\widehat{f}$, making it convenient for processing optimization. Different considerations that should be taken into account when formulating $\alpha(\mathbf{x})$ from $\widehat{f}$ include, for example, the value of the objective function [14], the information about the minimum location [15, 16, 17], and the information about the minimum value [18].
@@
Various algorithms have been developed to address the classic exploitation–exploration dilemma in BO, where exploitation involves selecting new solutions anticipated to improve the objective function value immediately and exploration focuses on sampling new solutions to reduce uncertainty in predictions of the objective function for a long-term improvement in the solution. An effective BO algorithm is deemed to balance these opposing concerns [4]. Notable deterministic acquisition functions, such as expected improvement (EI) [19], weighted EI [20], GP upper confidence bound (GP-UCB) [21], knowledge gradient (KG) [22], and likelihood-weighting [23], can manage the exploitation–exploration trade-off. However, they are considered myopic policies, and therefore tend to favor exploitation [14].
@@
Thompson sampling (TS) is a preferred algorithm to address the exploitation–exploration dilemma in solving multi-armed bandit problems [24, 25, 26, 27]. More specifically, TS selects arms (or actions) from a finite set of possible arms over a limited number of iterations. Each arm corresponds to a stochastic reward drawn from an unknown distribution. The goal is to craft a sequence of arms that maximizes the cumulative reward assuming that the rewards are independent of time and conditionally independent given the selected arms. When applied to GP-based BO, TS (or GP-TS [4]) generates a sequence of new solution points using the mechanism that involves random sampling from unknown posterior distributions of the global minimum location $\bf{x}^{\star}$ [28], which is due to the imperfect knowledge of $f$ described by the probabilistic model $\widehat{f}$. Rather than finding the posterior distributions directly, TS generates functions from $\widehat{f}$ and minimizes them for samples of $\bf{x}^{\star}$. Thus, we can consider TS a BO method that minimizes stochastic acquisition functions generated from the model posteriors for selecting new solution points. It is also worth noting that several information-theoretic optimization policies of BO, such as predictive entropy [17] and max-value entropy [18], compute their acquisition functions based on a set of samples generated by TS.
@@
While TS naturally manages the exploitation–exploration trade-off, its randomness can reduce the role of exploitation in optimization. When the GP posterior exhibits high uncertainty at the beginning of the optimization process, TS prioritizes exploration as it can diversify the selection of new solutions when limited information has been gained about the optimal solution $\bf{x}^{\star}$. As the number of observations increases and the GP posterior becomes concentrated, the algorithm transitions to exploiting knowledge about the true objective function. This exploitation strategy, however, is inferior due to the randomness of TS, motivating the quest for an intervention to improve its exploitation.
@@
Several works have explored the $\varepsilon$-greedy policy in BO and multi-armed bandit problems. De Ath et al. [30] proposed two schemes for applying the policy to BO. The first scheme performs exploration (with probability $\varepsilon$) by randomly selecting a point on the Pareto frontier obtained by simultaneously minimizing the posterior mean and maximizing the posterior standard deviation. The second scheme performs exploration by randomly selecting a point in the input variable space. Both schemes implement exploitation (with probability $1-\varepsilon$) by minimizing the posterior mean function. Jin et al. [31] introduced the so-called $\varepsilon$-exploring TS ($\varepsilon$-TS) to multi-armed bandit problems. Given the posterior distributions of arms, exploration (with probability $\varepsilon$) sets the estimated reward function for each arm as a random sample drawn from the associated posterior distribution, while exploitation (with probability $1-\varepsilon$) sets the estimated reward function as the sample mean function. Yet the performance of $\varepsilon$-greedy TS for BO remains unexplored.
Our contributions are as follows: (1) $\varepsilon$-greedy TS to BO, which is a simple, effective method to improve the exploitation of TS; and (2) empirical evaluations demonstrating that $\varepsilon$-greedy TS with an appropriate $\varepsilon$ is more robust than its two extremes (i.e., the generic TS and the sample-average TS), matching or outperforming the better of the two across various problems.
@@ 2.1 Overview of Gaussian Processes
Let $\mathcal{D}=\left\{\left({\bf X},{\bf y}\right)\right\}=\left\{\left({\bf x}^{i},y^{i}\right)\right\}_{i=1}^{N}$ denote a training dataset, where $\textbf{x}^{i}$ are observations of a $d$-dimensional vector of input variables and $y^{i}$ the corresponding observations of the objective function. We wish to build from $\mathcal{D}$ an observation model $y({\bf x})=f({\bf x})+\varepsilon_{\text{n}}:\mathbb{R}^{d}\mapsto\mathbb{R}$, where $\varepsilon_{\text{n}}\sim\mathcal{N}(0,\sigma^{2}_{\text{n}})$ is additive zero-mean Gaussian noise with variance $\sigma^{2}_{\text{n}}$. This observation noise is assumed to be independent and identically distributed.
A GP model assumes that any finite subset of an infinite set of objective function values has a joint Gaussian distribution [32]. Oftentimes this assumption is encoded using the following GP prior:
@@
where $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})=\text{cov}[f({\bf x}),f({\bf x}^{\prime})]:\mathbb{R}^{d}\times\mathbb{R}^{d}\mapsto\mathbb{R}$ is a positive semi-definite covariance function parameterized by a vector $\boldsymbol{\theta}$ of hyperparameters. Common-used covariance functions include squared exponential (SE), Matérn 3/2, Matérn 5/2, and automatic relevance determination squared exponential (ARD SE) [4, 32]. By conditioning on $\mathcal{D}$ the model prior given in Eq. 2 and utilizing the observation noise assumption, it is straightforward to show that the vector of observations $\left[y({\bf x}^{1}),\dots,y({\bf x}^{N})\right]^{\intercal}$ is distributed according to an $N$-variate Gaussian with zero mean and covariance matrix ${\bf C}={\bf K}+\sigma^{2}_{\text{n}}{\bf I}_{N}$, where the $(i,j)$th element of ${\bf K}$ is $\kappa({\bf x}^{i},{\bf x}^{j}|\boldsymbol{\theta})$ and ${\bf I}_{N}$ the $N$-by-$N$ identity matrix. By further applying the conditional multivariate Gaussian [33], we obtain the posterior predictive distribution of the objective function value at an unseen input variable vector ${\bf x}_{\star}$, i.e., $p\left(f({\bf x}_{\star})|{\bf x}_{\star},\mathcal{D}\right)=\mathcal{N}\left(\mu_{\text{f}}({\bf x}_{\star}),\sigma_{\text{f}}^{2}({\bf x}_{\star})\right)$, which encodes information about the model we wish to build. The mean and variance of the predictive distribution read
@@
Training GP models involves finding an optimal set of hyperparameters $\boldsymbol{\theta}$ that ensures robust predictive performance of trained GPs across unseen input variables. Exact training methods emulating all possible hyperparameter hypotheses are often impractical, leading to the prevalent use of approximate training methods. These approximate training methods fall into two main categories: deterministic approximations and Monte Carlo methods [34]. Deterministic approximations include the maximum likelihood method identifying a set of hyperparameters that maximizes the parameter likelihood $p(\mathcal{D}|\boldsymbol{\theta})$, and Laplace’s method approximating the posterior $p(\boldsymbol{\theta}|\mathcal{D})$ as a multivariate Gaussian with a posterior mean equal to the maximum likelihood estimation and a posterior covariance estimated from Laplace approximation. Monte-Carlo methods, such as importance sampling, the Metropolis, Gibbs sampling, and slice sampling, generate approximate samples of $\boldsymbol{\theta}$ from the posterior $p(\boldsymbol{\theta}|\mathcal{D})$ given the likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ and a prior $p(\boldsymbol{\theta})$. Several of the aforementioned approximate training methods have been incorporated in reliable GP toolboxes such as DACE [35], GPML [36], GPstuff [37], pyGPs [38], and GPflow [39].
@@
1:input: input variable domain $\mathcal{X}$, number of initial observations $N$, threshold for number of BO iterations $K$, standard deviation of observation noise $\sigma_{\text{n}}$
@@ 2.2 Bayesian Optimization via Thompson Sampling
Figure 1 shows sample paths generated from the GP posterior of univariate function $f(x)=x\sin{x}$ for $x\in[0,20]$ and an approximate distribution derived from their minimum locations. Specifically, we build from ten observations a GP model with zero mean and SE covariance function. From this GP model, we generate 50 sample paths, and five of them are shown in Fig. 1(a). Minimizing the generated sample paths and utilizing a kernel density estimation on the obtained solutions result in an approximate distribution $p({\bf x}^{\star}|\mathcal{D})$, as shown in Fig. 1(b). Note that the generic TS in Algorithm 1 generates and minimizes only one sample path in each iteration and does not attempt to approximate the posterior distribution of the minimum location. The following section describes how we can randomly generate sample path $g({\bf x}|\mathcal{D}^{k-1})$ in 10 of Algorithm 1.
2.3 Sampling from Gaussian Process Posteriors
Table 1: Spectral density functions for common-used stationary covariance functions [40].
Covariance $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ Spectral density function $S({\bf s})$
@@ 2.3 Sampling from Gaussian Process Posteriors
Given the GP posterior $\widehat{f}^{k}({\bf x})|\mathcal{D}^{k-1}$ characterized by a stationary covariance function, we follow the spectral sampling approach to generate a sample path $g({\bf x}|\mathcal{D}^{k-1})$, see e.g., [41] and [17]. This approach approximates the GP prior in Eq. 2 using a Bayesian linear model of randomized basis functions to avoid the computational cost due to exhaustive sampling from marginal distributions of the objective function values at finite sets of input locations, which scales cubically in the number of input locations. Such an approximation is rooted in Bochner’s theorem that guarantees the existence of a Fourier dual $S({\bf s})$ (${\bf s}\in\mathbb{R}^{d}$) of the stationary covariance function, which is called spectral density when a finite non-negative Borel measure is interpreted as a distribution [42]. The spectral density functions associated with several covariance functions of Matérn class and ARD SE are listed in Table 1. Once $S({\bf s})$ is determined, we can represent $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ using the following randomized feature map [41]:
@@
where the feature map $\boldsymbol{\phi}({\bf x})\in\mathbb{R}^{N_{\text{p}}}$ can be approximated by [41, 17]
@@
Here ${\bf W}\in\mathbb{R}^{N_{\text{p}}\times d}$ and ${\bf b}\in\mathbb{R}^{N_{\text{p}}}$ stack $N_{\text{p}}$ spectral points generated from the normalized spectral density $p({\bf s})=S({\bf s})/\kappa(0|\boldsymbol{\theta})$ and $N_{\text{p}}$ points drawn from the uniform distribution $\mathcal{U}[0,2\pi]$, respectively. $\kappa(0|\boldsymbol{\theta})$ is well defined because $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})$ is a stationary covariance function satisfying $\kappa({\bf x},{\bf x}^{\prime}|\boldsymbol{\theta})=\kappa(\left\|{\bf x}-{\bf x}^{\prime}\right\||\boldsymbol{\theta})$.
@@
1:Input: dataset $\mathcal{D}^{k-1}$, type of stationary covariance function $\kappa(\cdot,\cdot|\boldsymbol{\theta})$, number of spectral points $N_{\text{p}}$, standard deviation of observation noise $\sigma_{\text{n}}$
@@
Figure 2 shows approximate SE and Matérn 5/2 covariance functions computed from Eq. 6 for different numbers of spectral point samples. We see that the approximate covariance functions converge to the true ones when the number of spectral points increases.
where the prior of $\boldsymbol{\beta}$ is $\mathcal{N}({\bf 0},{\bf I}_{N_{\text{p}}})$. By further conditioning Eq. 8 on the data, we obtain the following mean and covariance of the posterior of $\boldsymbol{\beta}$ [17]
where $\boldsymbol{\Phi}=[\boldsymbol{\phi}({\bf x}),\dots,\boldsymbol{\phi}({\bf x}^{N})]^{\intercal}\in\mathbb{R}^{N\times N_{\text{p}}}$. To this end, we can approximate the GP posterior using a Bayesian linear model that weights the randomized basis functions using samples from the posterior of $\boldsymbol{\beta}$.
The generation of $g({\bf x}|\mathcal{D}^{k-1})$ using random features is detailed in Algorithm 2. Note that 10 of Algorithm 2 involves the Cholesky decomposition of matrix $\boldsymbol{\Phi}^{\intercal}\boldsymbol{\Phi}+\sigma_{\text{n}}^{2}{\bf I}_{N_{\text{p}}}$ to facilitate the calculation of its inverse. Alternative posterior sampling techniques such as the decoupled method [43], which generates GP posteriors by updating the corresponding GP prior samples, can be adopted to increase the sample path accuracy.
3 Epsilon-greedy Thompson Sampling
In this section, we first introduce another version of TS called sample-average TS (Section 3.1). We then describe $\varepsilon$-greedy TS using the sample-average and generic TS in the context of $\varepsilon$-greedy policy (Section 3.2).
3.1 Sample-average Thompson Sampling
Another approach to generate an average sample path is to use a decoupled representation via pathwise conditioning [45, 46], which can be computed at the same cost of generating one sample path, regardless of the nominal number $N_{s}$ of sample paths being averaged.
The sample-average TS and the generic TS can be considered two extremes of TS. If $N_{\text{s}}=\infty$, $g({\bf x}|\mathcal{D}^{k-1})$ in Eq. 10 is indeed the GP posterior mean in Eq. 3 whose minimum promotes exploitation. If $N_{s}=1$, the sample-average TS recovers the generic TS that favors exploration. Thus, there exists an optimal state between these two extremes that corresponds to an unknown value of $N_{s}$ where TS balances exploitation and exploration. However, we do not attempt to find such an optimal state in this work. We instead set $N_{\text{s}}$ at a sufficiently large value, say $N_{\text{s}}=50$, to enforce the exploitation of TS.
3.2 $\varepsilon$-greedy Thompson Sampling
The two distinct extremes of TS motivate the use of $\varepsilon$-greedy policy to randomly switch between them. Specifically, we implement the generic TS with probability $\varepsilon$ to explore the input variable space. We invoke the sample-average TS with probability $1-\varepsilon$ to guide the search toward exploitation. We do not use the posterior mean in Eq. 3 for exploitation because we observe that computing its derivatives is more expensive than computing the derivatives of the average sample path in Eq. 10, which adds computational cost if optimization is performed via a gradient-based algorithm. It is also straightforward to recover the generic TS from $\varepsilon$-greedy TS by simply setting $\varepsilon=1$ or $N_{\text{s}}=1$. Algorithm 3 details the proposed $\varepsilon$-greedy TS.